OpenStudy (anonymous):

what is 0.783 as a fraction in simplest form

OpenStudy (anonymous):

0.783 as a fraction is 783/100

OpenStudy (anonymous):

and 783/100 is as simplified as u can get

Directrix (directrix):

.78 = 78/100 .783 = 783 / 1000 Do 783 and 1000 have any common factors? If not, .783 will be 783 / 1000 1,3,9,27,29,87,261,783 are the factors of 783

Directrix (directrix):

@SeaRocks77 >> 0.783 as a fraction is 783/100 Should be .783 as a fraction is 783 / 1000